OpenStudy (anonymous):

help please ? D; I dont understand this . - You can cut a pizza into two pieces with one straight cut. With two cuts you can get four pieces. Three cuts give a maximum of seven pieces. Extend the pattern (either by drawing or creating a table) to determine what is the maximum number of pieces with four cuts. a. 11 pieces c. 16 pieces b. 15 pieces d. 21 pieces

OpenStudy (kewlgeek555):

Okay. This the pattern: 1 cut = 2 pieces 2 cuts = 4 pieces 3 cuts = 7 pieces Okay, it is confusing but I found it out. 2 + 2 = 4 4 + 3 = 7 So now do 7 + 4 = your answer!

OpenStudy (anonymous):

ohh well when you put i like that it seems so easy ! thanks!

OpenStudy (kewlgeek555):

You are very welcome.